# In this paper, we consider the problem whether there exist transcendental meromorphic functions which satisfy the functional equation

### v4

@@ 1. Introduction and main results
satisfying as where is the Nevanlinna characteristic of Following [17, p.466], we define the notion of the small function of a non-constant meromorphic function in the complex plane as follows: let be a non-constant meromorphic function. and let be also a meromorphic function in the complex plane such that for all possibly outside a set of finite linear measure. Then, the meromorphic function is called small compared to or slowly moving with respect to or a small function of for short.
@@
For the existence of non-constant meromorphic solutions of a much more general three term functional equation, we recall the following result from [20] and Yang-Zhang [37, 38]:
@@ 2. Preliminaries
then uniformly in some neighbourhood of every point of and hence by the Heine-Borel theorem uniformly on every compact subset of see, for example, [18, pp.157-160]. We recall the following classical normality criterion of a class of meromorphic functions due to Marty[25]:
@@
(Pang-Zalcmanβs Lemma, [31, 43]).
@@
(iii) β functions such that spherically uniformly on compact subset of
where is a non-constant meromorphic function. The function
@@ 3. Notations and definitions for the proofs
([23, Definition 1]) Let be a positive integer and Next we denote by the counting function of those -points of (counted with proper multiplicities) whose multiplicities are not greater than and denote by the counting function of those -points of (counted with proper multiplicities) whose multiplicities are not less than We denote by and the reduced forms of and respectively. Here and
@@
the reduced counting function of common zeros of in Restrictions can be given in these definitions. For instance, the notation will mean the reduced counting function of common zeros of and in that are not zeros of The notation will mean the reduced counting function of common zeros of and in that are not zeros of and In addition, the notation
denotes the counting function of those zeros of in that are zeros of where each point in
is counted according to its
multiplicity as a zero of The notation
denotes the counting function of those zeros of in that are poles of where each point in
is counted according to its multiplicity as a zero of We denote by and the reduced forms of and respectively.
The notation denotes the reduced counting function of those simple zeros of in that are zeros of
The notation denotes the reduced counting function of those simple zeros of in that are poles of
Thirdly, we denote by the reduced counting function of common simple poles of and in denote by the reduced counting function of common simple zeros of and in denote by the reduced counting function of those zeros of in that are not zeros of and denote by the reduced counting function of those zeros of in that are neither zeros of nor zeros of
Last, we denote by the reduced counting function of those common simple zeros in where each such common simple zero of and is neither a zero of nor a zero of
@@ 4. Proof of Theorem 8
for each positive integer Here is some small positive number such that is analytic in for each positive integer and such that
@@
and with are analytic functions on and is the only one zero of in by Hurwitzβs theorem (cf.[32, p.9 Hurwitz Theorem]) we deduce that has one zero in at least for the large positive integer However, from (99), (100), the left inequality of (98) and the supposition we deduce that
@@
Combining this with the known result that spherically uniformly on the closed disk where and with are analytic functions on and is the only one zero of in we deduce by Hurwitzβs theorem (cf.[32, p.9 Hurwitz Theorem]) that has one zero in at least for the large positive integer However, from (99), (100), the left inequality of (98) and the supposition we deduce that
@@
for the large positive integer we deduce by (102) and Weierstrass theorem (cf.[32, p.9, Weierstrass Theorem]) that
@@
is analytic on the closed disk such that for each and is the only one zero of on the closed disk we deduce by (103), Hurwitzβs theorem (cf.[32, p.9 Hurwitz Theorem]) and the supposition that the composition is analytic on the closed disk
@@ 5. Proof of Theorem 9
From Lemma 15 and the supposition that and are transcendental entire functions, we deduce that has infinitely many repelling periodic points of minimal period for the given positive integer such that As in the proof of Theorem 8, next we let be a repelling periodic point of minimal period and we have from Lemma 16 that As in the proof of Theorem 8, we have (88)-(98). Noting that we deduce by (90), (91) and the definition of the Julia set of the transcendental entire function that the family with for each positive integer is not normal at the point Accordingly, it follows from (96) and (97) that the family defined in is not normal at the origin point Combining this with Lemma 14, we see that there exist (iv) points (v) positive numbers

### v5

@@ 1. Introduction and main results
satisfying as where is the Nevanlinna characteristic of Following [18, p.466], we define the notion of the small function of a non-constant meromorphic function in the complex plane as follows: let be a non-constant meromorphic function. and let be also a meromorphic function in the complex plane such that for all possibly outside a set of finite linear measure. Then, the meromorphic function is called small compared to or slowly moving with respect to or a small function of for short.
@@
For the existence of non-constant meromorphic solutions of a much more general three term functional equation, we recall the following result from Hu-Li-Yang [21] and Yang-Zhang [39, 40]:
@@ 2. Preliminaries
then uniformly in some neighbourhood of every point of and hence by the Heine-Borel theorem uniformly on every compact subset of see, for example, [19, pp.157-160]. In addition, we need the following notion of the spherical derivative of a meromorphic function in the complex plane (cf.[19, 38]: let be a non-constant meromorphic function. The spherical derivative of
at is given as We recall the following classical normality criterion of a class of meromorphic functions due to Marty[26]:
@@
The following result is due to Pang [33]:
Lemma 14.
([33, Lemma 1])
Let be a meromorphic function in the domain and let be a real number satisfying If there exists a point where is a fixed positive number satisfying such that
then there exist a point and a real number satisfying such that
@@
(Pang-Zalcmanβs Lemma, [17, 32, 45]).
@@
there exist, for each
(i) points
(ii) positive numbers
and
(iii) functions such that spherically uniformly on compact subset of
where is a non-constant meromorphic function. The function
may be taken to satisfy the normalization
From Remark 2 (III) and the proof of Lemma 15 for in Remark 2 (II), we get the following result:
Lemma 16.
Let be a family of holomorphic functions in a disc where is a point, is a domain in the complex plane, and is an infinite sequence of positive numbers such that for each positive integer and let be a real number
satisfying Then, if is not normal at the point
@@
(iii) β an infinite subsequence of the family say itself, such that spherically uniformly on compact subset of
where is a non-constant entire function. The function
@@ 3. Notations and definitions for the proofs
([24, Definition 1]) Let be a positive integer and Next we denote by the counting function of those -points of (counted with proper multiplicities) whose multiplicities are not greater than and denote by the counting function of those -points of (counted with proper multiplicities) whose multiplicities are not less than We denote by and the reduced forms of and respectively. Here and
@@
the reduced counting function of common zeros of in Restrictions can be given in these definitions. For instance, the notation will mean the reduced counting function of common zeros of and in that are not zeros of The notation will mean the reduced counting function of common zeros of and in that are not zeros of and The notation will mean the reduced counting function of common zeros of in that are either zeros of or zeros of
Thirdly, we denote by the reduced counting function of common simple poles of and in and denote by the reduced counting function of common simple zeros of and in
@@ 4. Proof of Theorem 8
for each positive integer Here is some small positive number satisfying for each positive integer such that is analytic in for each positive integer and such that
@@
and with are analytic functions on and is the only one zero of in by Hurwitzβs theorem (cf.[34, p.9 Hurwitz Theorem]) we deduce that has one zero in at least for the large positive integer However, from (123), (124), the left inequality of (122) and the supposition we deduce that
@@
Combining this with the known result that spherically uniformly on the closed disk where and with are analytic functions on and is the only one zero of in we deduce by Hurwitzβs theorem (cf.[34, p.9 Hurwitz Theorem]) that has one zero in at least for the large positive integer However, from (123), (124), the left inequality of (122) and the supposition we deduce that
@@
for the large positive integer we deduce by (126) and Weierstrass theorem (cf.[34, p.9, Weierstrass Theorem]) that
@@
is analytic on the closed disk such that for each and is the only one zero of on the closed disk we deduce by (127), Hurwitzβs theorem (cf.[34, p.9 Hurwitz Theorem]) and the supposition that the composition is analytic on the closed disk
@@ 5. Proof of Theorem 9
From Lemma 17 and the supposition that and are transcendental entire functions, we deduce that has infinitely many repelling periodic points of minimal period for the given positive integer such that As in the proof of Theorem 8, next we let be a repelling periodic point of minimal period and we have from Lemma 18 that As in the proof of Theorem 8, we have (112)-(122), where and in what follows, is some small positive number satisfying for each positive integer Noting that we deduce by (114), (115) and the definition of the Julia set of the transcendental entire function that the family with for each positive integer is not normal at the point Accordingly, it follows from (120) and (121) that the family defined in is not normal at the origin point Combining this with Lemma 16, we see that there exist (iv) points in such that (v) positive numbers such that